OpenStudy (anonymous):

what is the solution set of 4m=-84 do i divide

OpenStudy (anonymous):

x=7. (add 4 to both sides: x=3+4=7)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what is the solution set of -8+a=5

OpenStudy (anonymous):

13